OpenStudy (anonymous):

4(2x-1/2)-3(x-1/3)=2

OpenStudy (saifoo.khan):

find x?

OpenStudy (anonymous):

solve the following equations

OpenStudy (saifoo.khan):

x =1

OpenStudy (anonymous):

join my online classroom and I'll help http://www.twiddla.com/562616

OpenStudy (anonymous):

unless you just wanted the answer

OpenStudy (anonymous):

the answers 3/5 show me the steps

OpenStudy (anonymous):

I sent you a link to my classroom, you haven't clicked it yet.

jhonyy9 (jhonyy9):

so newly 4(2x-1/2)-3(x-1/3)=2 --- 2(4x-2)/2-3(x-1)/3=2 --- 4x-2-x+1=2 3x=3 --- x=3/3 =1